OpenStudy (konradzuse):

integral of 3/(x^2+9) from sqrt(3) to 3

OpenStudy (konradzuse):

\[\int\limits_{\sqrt{3}}^{3} \frac{3}{x^2 + 9}\]

OpenStudy (anonymous):

pull out the 3 on top and you'll see what you have

OpenStudy (anonymous):

also you can switch the x^2+9 to 9+x^2

OpenStudy (konradzuse):

yup it's arctan

OpenStudy (konradzuse):

I'm just trying to find out how to get the answer of pi/12.

OpenStudy (konradzuse):

\[3\int\limits_{\sqrt{3}}^{3} \frac{1} {x^2 +9} = \frac{1}{3} \arctan \frac{x}{3}\]

OpenStudy (konradzuse):

actually I think I know what I did, I replaced the 1 with sqrt{3} and 3 lol.

OpenStudy (konradzuse):

I was wondering why it made no sense :P